# Injectives obstruct Fourier-Mukai functors

### v1

@@ 1. Introduction
The idea behind Fourier-Mukai functors as categorified correspondences was introduced by S. Mukai [Muk81] to understand the relation between the derived category of an abelian variety $\mathcal{A}$ and its dual $\mathcal{A}^{\vee}$. Its classical definition being:
@@
for two schemes $X$ and $Y$ is Fourier-Mukai if there exists an $M$ in $\mathcal{D}^{b}\left(\mathrm{coh}\left(X\times Y\right)\right)$, respectively in $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\times Y\right)\right)$, such that
@@
This concept allows the study of functors from a geometric point of view and in particular the application of geometric arguments and constructions to them. With time it was realized that in fact every known functor between derived categories of smooth projective schemes is Fourier-Mukai. For example push-forward and pull-back along a morphism of schemes $f$ is induced by taking $M$ to be the structure sheaf of its graph $\mathcal{O}_{\Gamma_{f}}$.
A celebrated result by D. Orlov gave a criterion to check whether a functor is Fourier-Mukai,
@@
[Orl97] Let $f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$ be a fully-faithful functor between derived categories of smooth projective schemes. Then $f$ is Fourier-Mukai.
Whereas B. Toën realized a beautiful conceptual categorization of the property,
@@
where $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ is a $\mathrm{dg}$-enhancement of $\mathcal{D}^{b}\left(\_\right)$.
This Theorem in particular allows the interpretation of Fourier-Mukai functors as the geometric analog to bimodules and their represented functors. These observations led to a conjecture that actually all exact functors between derived categories of smooth projective schemes are Fourier-Mukai.
@@
We were able to generalize their argument to higher dimensions in [Kü22] using twisted Hodge diamonds, while T. Raedschaelders, A. Rizzardo and M. Van den Bergh [RRVdB19] were able to construct similar non-Fourier-Mukai functors by losing control over the target space.
A big drawback of all these approaches is that they heavily rely on the existence of tilting objects in the source category in order to control liftability. We avoid this issue here by replacing the tilting bundle with the subcategory of injectives in the canonical heart of the derived category of quasi-coherent sheaves. This observation allows the computation of the equivariant characteristic morphism with respect to the category of injectives and explicit computations of the induced $\mathcal{A}_{\infty}$-obstructions. In particular we are able to prove the following very general result:
@@
Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface. Then the space of non-Fourier-Mukai functors arising as deformation of the shape $\psi_{\eta}$ can be computed explicitly.
@@ 1.1. Proof strategy
The key insight in our approach is that the non-liftability of $\psi_{\eta}\left(T\right)$ boils down to the the functor $\psi_{\eta}\circ\iota_{T}:*_{\Gamma}\to\mathcal{D}^{b}\left(Y\right)$ not admitting an $\mathcal{A}_{\infty}$-lift, here we denoted by $*_{\Gamma}$ the $𝕜\Bbbk$-linear category with one object * and endomorphisms $\Gamma$. In particular, since $T$ was assumed to be tilting the functor
@@
In order to explain our proof strategry, consider the follwoing diagram
@@
As $\iota$ uniquely determines $\mathrm{dg}$-functors from the category $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\right)\right)$ we can compute its equivariant characteristic morphism in order to obstruct an $\mathcal{A}_{\infty}$-lift of $L$. This yields a non-vanishing $\mathcal{A}_{\infty}$-obstruction $o_{m}\left(L\circ\iota\right)$. The plan is to push this obstruction forward. In contrast to [RVdBN19] and [Kü22] this needs to be done with a lot of care as we do not have access to Serre duality in our setting. In particular we study the concrete construction of $\widetilde{f}_{*}$ to compute that the obstruction $o_{m}\left(\psi_{\eta}\circ\iota\right)$ indeed is independent of choices and coincides with the pushed forward obstruction $f_{*}o_{m}\left(L\circ\iota\right)$. This means that there cannot be an $\mathcal{A}_{\infty}$-lift of $\psi_{\eta}\circ\iota$. In particular, as $\iota$ is the inclusion of injectives, it is itself a $\mathrm{dg}$-functor and hence we conclude that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift as desired.
@@ 1.2. Structure of the paper
In Section 2.1 we recall preliminaries on Hochschild coholomology and the characteristic morphism in the classical setting. In particular we recall in Definition 2.5 the construction of an infinitesimal deformation along a Hochschild cohomology class, which is a key player in the construction of $\psi_{\eta}.$
@@
Since we are not using tilting objects in the category of coherent sheaves we are not able to use Serre duality as in [RVdBN19, Kü22] in order to show that the relevant $\mathrm{Ext}$-groups are concentrated in degree $-1$. So we will have to control the $\mathcal{A}_{\infty}$-structure on our target category more explicitly. For this we study in Section 6 the explicit structure of the functor $\widetilde{f}$. In particular we completely understand the essential image and its $\mathcal{A}_{\infty}$-bimodule structure.
Then we finish in Section 7 by computing explicitly that the obstructions of $\psi_{\eta}$ cannot vanish. We then combine this result with computations from [Kü22] in order to conclude the existence of a vast amount of non-Fourier-Mukai functors with well-behaved target categories.
@@ Acknowledgements
I would like to thank the SL-Math institute for hosting the semester on “non-commutative algebraic geometry” where a lot of this work has been completed. Furthermore I want to show my gratitude to M. Van den Bergh and A. Rizzardo for helpful discussions and deep insights and Š. Špenko for her awesome support and helpful discussions. Finally I would like to thank B. Briggs and A. Cǎldǎraru for their insights on Hochschild cohomology, $\mathcal{A}_{\infty}$-structures and formality of sheaves over source spaces.
@@
Let $𝕜\Bbbk$ denote a field of characteristic zero. We consider all schemes and categories to be over $𝕜\Bbbk$. For a $𝕜\Bbbk$-linear category $\mathcal{X}$, denote by $\mathcal{D}^{b}\left(\mathcal{X}\right)$ the bounded derived category of modules over $\mathcal{X}$. We will consider left modules. Denote by $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ the (canonical) dg-enhancement of $\mathcal{D}^{b}\left(\_\right)$. We denote by $\mathcal{D}^{b}_{\mathcal{C}}\left(\_\right)$ the full subcategory of $\mathcal{D}^{b}\left(\_\right)$ with cohomology objects in $\mathcal{C}$. We will also refer to $\mathrm{dg}$ respectively $\mathcal{A}_{\infty}$-functors between smooth geometric categories as Fourier Mukai functors following the insight from B. Toën [Toe07]. We also use the shorthand
@@
for a morphism $\mathcal{A}^{\otimes i}\to\mathcal{M}$ and $\mathcal{M}$ an $\mathcal{A}_{\infty}$-$\mathcal{A}$-module.
@@ 2. Preliminaries
We start off by recalling some basic notions needed for our constructions and arguments. For the basic notions on $\mathcal{A}_{\infty}$-structures we refer to the notes by B. Keller [Kel01], the work of H. Lefevre [LH03] and P. Seidel [Sei08]. For the construction of $\mathrm{dg}$-hulls we refer to [KKS09].
2.1. Preliminaries Hochschild cohomology and characteristic morphism
As we will work a lot with obstructions and Hochschild cohomology we recall the basic constructions and notions here. For a more in depth discussion on the incarnations of Hochschild cohomology we use here, we refer the reader to the original work of G. Hochschild [Hoc45], R. Buchweitz and H. Flenner [BF08] respectively R. Swan [Swa96] for its geometric incarnation, and for its relation to deformation theory to the work of M. Van den Bergh and W. Lowen [LVdB05]. Finally we want to refer for a treatment on the characteristic morphism to the work of B. Briggs and V. Gelinas [BG17].
@@ 2.1. Preliminaries Hochschild cohomology and characteristic morphism
$df\left(x_{1}\otimes\cdot\otimes x_{n+1}\right):=$ $x_{1}f\left(x_{2}\otimes\cdots\otimes x_{n+1}\right)$
$+\sum_{i=1}^{n}\left(-1\right)^{i}f\left(x_{1}\otimes\cdots\otimes x_{1}x_{i+1}\otimes\cdots\otimes x_{n+1}\right)$
@@
The Hochschild cohomology $\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)$ is the cohomology of $\mathrm{CC}^{*}\left(\mathcal{X},\mathcal{M}\right)$.
The following classical Lemma 2.2 allows for a more streamlined construction and will help us interpret the Hochschild cohomology in more generality later. In particular it highlights the importance of $\mathcal{M}$ being a bimodule.
@@
In bigger generality constructing tensor products of $\mathcal{A}_{\infty}$-categories needs to be done with a lot of care [LRG20]. However, we will only consider tensor products with $𝕜\Bbbk$-linear categories, which are a lot better behaved and allow our naive approach to work out.
The restriction in the following definition to $m\geq 3$ is not necessary in general. However, as we will have naturally this assumption and it makes the formulas a lot more concise we present it in this special case.
@@
$\left(x,m\right)\otimes\left(x^{\prime},m^{\prime}\right)$ $\mapsto\left(xx^{\prime},xm+m^{\prime}x\right)$
@@
$x_{1}\otimes...\otimes x_{m}$ $\mapsto\eta\left(x_{1},...,x_{m}\right)$
@@ 2.2. Equivariant objects over categories
In this section we recall the basic notions of equivariant objects and how to interpret them as functors from a small category, which we will apply in the next section to compute the equivariant characteristic morphisms of these objects.
@@
There also exists a $\Gamma$-equivariant incarnation of the characteristic morphism, given by:
@@
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category, $\mathcal{M}$ a $𝕜\Bbbk$-central bimodule and let $\mathcal{G}$ be a $\Gamma$-equivariant $\mathcal{X}$-module. Then _$\_\otimes\mathcal{G}$ induces the the $\Gamma$-equivariant characteristic morphism
@@
While $\mathrm{RHom}\left(\_,\mathcal{G}\right)$ induces the the $\Gamma$-equivariant dual characteristic morphism
@@
We can do the same construction the other way around, i.e. make a $\Gamma$-equivariant object $X$ into a functor via fixing the image of * to be $X$ and the structure morphism the map on morphism spaces.
@@
We use the following notation for the category of equivariant objects over $I$ in $\mathcal{C}$ by
@@ 3. Equivariant characteristic morphism of Injectives
We study the characteristic morphism introduced in the previous chapter for the subcategory of injectives in a category with enough injectives. We start by discussing how we can reinterpret the characteristic morphism by the work of [Toe07]. We then use this insight to prove that we can treat the full subcategory of injectives similarly to a tilting object.
@@ 3.1. Interpreting characteristic morphisms of a subcategory as pullback
We use observations from [Toe07] to study in more depth the characteristic morphism
@@
$c_{\Gamma,G}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\Gamma}}^{*}\left(G,\mathcal{M}\otimes_{\mathcal{C}}G\right).$
@@
Let $\mathcal{C}$ be an abelian category and $I\subset\mathcal{C}$ a subcategory. Then denote by $\iota_{I}:I\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{C}\right)$ the inclusion of objects of $I$.
@@
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\Gamma,\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota_{\mathcal{G}},\psi_{\mathcal{M}}\circ\iota_{\mathcal{G}}\right)\cong\mathrm{Ext}_{\mathcal{C}_{\Gamma}}\left(\mathcal{G},\mathcal{G}\otimes\mathcal{M}\right).$$c_{\Gamma,\mathcal{G}}\cong\iota_{\mathcal{G},*}$
In particular we can use this interpretation to compute the equivariant characteristic morphism of the subcategory $I\in\mathcal{C}$:
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}-\mathrm{Cat}\right)\left(\mathcal{I},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota,\psi_{M}\circ\iota\right)\cong\mathrm{Ext}^{*}_{\mathcal{C}_{\mathcal{I}}}\left(\iota,\psi_{M}\circ\iota\right).$$c_{\iota_{I},I}=\iota_{I,*}$
The above observation is in particular useful as explicit computations for bimodules between $𝕜\Bbbk$-linear categories, or, even more, $\mathcal{A}_{\infty}$/$\mathrm{dg}$-categories can be very unwieldy. However, the above observation allows us to compute the equivariant characteristic morphisms of the $I$-$\mathcal{C}$ bimodule $\iota_{I}$ in special cases.

### v2

@@ Abstract.
We use injectives as a big tilting object to obstruct liftability of exact functors to the $\mathrm{dg}$-level.
@@ 1. Introduction
The idea behind Fourier–Mukai functors as categorified correspondences was introduced by S. Mukai [Muk81] to understand the relation between the derived category of an abelian variety $\mathcal{A}$ and its dual $\mathcal{A}^{\vee}$. The classical definition is:
@@
for two schemes $X$ and $Y$ is Fourier-Mukai if there exists a sheaf $M$ in $\mathcal{D}^{b}\left(\mathrm{coh}\left(X\times Y\right)\right)$, respectively in $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\times Y\right)\right)$, such that
@@
This concept allows for the study of functors from a geometric point of view and in particular the application of geometric arguments and constructions to them. With time it was realized that in fact every known functor between derived categories of smooth projective schemes is Fourier-Mukai. For example, push-forward and pull-back along a morphism of schemes $f$ is induced by taking $M$ to be the structure sheaf of its graph $\mathcal{O}_{\Gamma_{f}}$.
A celebrated result by D. Orlov gave a criterion to check whether a functor is Fourier-Mukai:
@@
[Orl97] Let $f:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$ be a fullyfaithful functor between derived categories of smooth projective schemes. Then $f$ is Fourier-Mukai.
Whereas B. Toën realized a beautiful conceptual categorization of the property.
@@
such that $\mathrm{H}^{0}\left(\widehat{f}\right)\cong f$, where $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ is a $\mathrm{dg}$-enhancement of $\mathcal{D}^{b}\left(\_\right)$.
This theorem in particular allows the interpretation of Fourier-Mukai functors as the geometric analog to bimodules and their represented functors. These observations led to a conjecture that all exact functors between derived categories of smooth projective schemes are Fourier-Mukai.
@@
We were able to generalize their argument to higher dimensions in [Kü24] using twisted Hodge diamonds, while T. Raedschaelders, A. Rizzardo and M. Van den Bergh [RRVdB19] were able to construct similar non-Fourier-Mukai functors by losing control over the target space.
A big drawback of all these approaches is that they heavily rely on the existence of tilting objects in the source category in order to control liftability, which tend to be rare. We avoid this issue here by replacing the tilting bundle with the subcategory of injectives in the canonical heart of the derived category of quasi-coherent sheaves. This observation allows the computation of the equivariant characteristic morphism with respect to the category of injectives and explicit computations of the induced $\mathcal{A}_{\infty}$-obstructions. In particular we are able to prove the following very general result:
@@
Let $f:X\hookrightarrow\mathbb{P}^{n+1}$ be a smooth degree $d$-hypersurface. Then the space of non-Fourier-Mukai functors arising as deformations of the shape $\psi_{\eta}$ can be computed explicitly.
Finally we show a possible avenue for generalizing our result to coherent sheaves by using extension results by J. Lurie [Lur09] respectively B. Toën [Toe07].
Theorem 8.12.
Assume that $\psi_{\eta}$ is continuous. Then the restriction
$\psi_{\eta}:\mathcal{D}^{b}\left(\mathrm{coh}X\right)\to\mathcal{D}^{b}\left(\mathrm{coh}Y\right)$
is not Fourier-Mukai.
@@ 1.1. Proof strategy
The key insight in our approach is that the non-liftability of $\psi_{\eta}\left(T\right)$ boils down to the the functor $\psi_{\eta}\circ\iota_{T}:*_{\Gamma}\to\mathcal{D}^{b}\left(Y\right)$ not admitting an $\mathcal{A}_{\infty}$-lift, here we denoted by $*_{\Gamma}$ the $𝕜\Bbbk$-linear category with one object * and endomorphisms $\Gamma$. In particular, since $T$ was assumed to be tilting, the functor
@@
In order to explain our proof strategy, consider the following diagram
@@
As $\iota$ uniquely determines $\mathrm{dg}$-functors from the category $\mathcal{D}^{b}\left(\mathrm{Qcoh}\left(X\right)\right)$ we can compute its equivariant characteristic morphism in order to obstruct an $\mathcal{A}_{\infty}$-lift of $L$. This yields a non-vanishing $\mathcal{A}_{\infty}$-obstruction $o_{m}\left(L\circ\iota\right)$. The plan is to push this obstruction forward. In contrast to [RVdBN19] and [Kü24] this needs to be done with a lot of care as we do not have access to Serre duality in our setting. In particular we study the concrete construction of $\widetilde{f}_{*}$ to compute that the obstruction $o_{m}\left(\psi_{\eta}\circ\iota\right)$ indeed is independent of choices and coincides with the pushed forward obstruction $f_{*}o_{m}\left(L\circ\iota\right)$. This means that there cannot be an $\mathcal{A}_{\infty}$-lift of $\psi_{\eta}\circ\iota$. In particular, as $\iota$ is the inclusion of injectives, it is itself a $\mathrm{dg}$-functor and hence we conclude that $\psi_{\eta}$ cannot have an $\mathcal{A}_{\infty}$-lift as desired.
@@ 1.2. Structure of the paper
In Section 2.1 we recall preliminaries on Hochschild cohomology and the characteristic morphism in the classical setting. In particular we recall in Definition 2.5 the construction of an infinitesimal deformation along a Hochschild cohomology class, which is a key player in the construction of $\psi_{\eta}.$
@@
Since we are not using tilting objects in the category of coherent sheaves we are not able to use Serre duality as in [RVdBN19, Kü24] in order to show that the relevant $\mathrm{Ext}$-groups are concentrated in degree $-1$. So we will have to control the $\mathcal{A}_{\infty}$-structure on our target category more explicitly. For this we study in Section 6 the explicit structure of the functor $\widetilde{f}$. In particular we completely understand its essential image and its $\mathcal{A}_{\infty}$-bimodule structure.
We finish in Section 7 by computing explicitly that the obstructions of $\psi_{\eta}$ cannot vanish. We then combine this result with computations from [Kü24] in order to conclude the existence of a vast amount of non-Fourier-Mukai functors with well-behaved target categories and showing possible arguments for the case of coherent sheaves.
@@ Acknowledgements
I would like to thank the SL-Math institute for hosting the semester on “non-commutative algebraic geometry” where a lot of this work has been completed. Furthermore I want to show my gratitude to M. Van den Bergh and A. Rizzardo for helpful discussions and deep insights and Š. Špenko for her awesome support and helpful discussions. I would like to thank B. Briggs and A. Cǎldǎraru for their insights on Hochschild cohomology, $\mathcal{A}_{\infty}$-structures and formality of sheaves over source spaces. Finally I are deeply grateful for the insight of J. Holstein on continuation of functors allowing the possible restriction to coherent sheaves.
@@
Let $𝕜\Bbbk$ denote a field of characteristic zero. We consider all schemes and categories to be over $𝕜\Bbbk$. For a $𝕜\Bbbk$-linear category $\mathcal{X}$, denote by $\mathcal{D}^{b}\left(\mathcal{X}\right)$ the bounded derived category of left modules over $\mathcal{X}$. For $X$ a scheme we denote by $\mathcal{D}^{b}\left(X\right)$ the bounded derived category of coherent sheaves. Denote by $\mathcal{D}_{\mathrm{dg}}^{b}\left(\_\right)$ the (canonical) dg-enhancement of $\mathcal{D}^{b}\left(\_\right)$. We denote by $\mathcal{D}^{b}_{\mathcal{C}}\left(\_\right)$ the full subcategory of $\mathcal{D}^{b}\left(\_\right)$ with cohomology objects in $\mathcal{C}$. We also use the slightly strange convention of $f_{*}$ for the pullback of $𝕜\Bbbk$-linear categories in order to mimic the geometric setting and improve readability. We will also refer to $\mathrm{dg}$ respectively $\mathcal{A}_{\infty}$-functors between smooth geometric categories as Fourier Mukai functors following the insight of B. Toën [Toe07]. Finally we use the shorthand
@@
for the Gerstenhaber bracket of a morphism $\mathcal{A}^{\otimes i}\to\mathcal{M}$ and $\mathcal{M}$ an $\mathcal{A}_{\infty}$-$\mathcal{A}$-module.
@@ 2. Preliminaries
We start by recalling some basic notions needed for our constructions and arguments. For the basics on $\mathcal{A}_{\infty}$-structures we refer to the notes by B. Keller [Kel01], the work of H. Lefevre [LH03] and P. Seidel [Sei08]. For the construction of $\mathrm{dg}$-hulls we refer to [KKS09].
2.1. Hochschild cohomology and characteristic morphism
As we will work a lot with obstructions in Hochschild cohomology we recall the basic constructions and notions here. For a more in depth discussion on the incarnations of Hochschild cohomology we use here, we refer the reader to the original work of G. Hochschild [Hoc45], R. Buchweitz and H. Flenner [BF08] respectively R. Swan [Swa96] for its geometric incarnation, and for its relation to deformation theory to the work of M. Van den Bergh and W. Lowen [LVdB05]. Finally we want to refer for a treatment on the characteristic morphism to the work of B. Briggs and V. Gelinas [BG17].
@@ 2.1. Hochschild cohomology and characteristic morphism
$df\left(x_{1}\otimes\cdots\otimes x_{n+1}\right):=$ $x_{1}f\left(x_{2}\otimes\cdots\otimes x_{n+1}\right)$
$+\sum_{i=1}^{n}\left(-1\right)^{i}f\left(x_{1}\otimes\cdots\otimes x_{i}x_{i+1}\otimes\cdots\otimes x_{n+1}\right)$
@@
The Hochschild cohomology $\mathrm{HH}^{*}\left(\mathcal{X},\mathcal{M}\right)$ is the cohomology of $\mathrm{CC}^{\cdot}\left(\mathcal{X},\mathcal{M}\right)$.
The following classical Lemma 2.2 allows for a more streamlined construction and will help us interpret the Hochschild cohomology in more generality later.
@@
In bigger generality constructing tensor products of $\mathcal{A}_{\infty}$-categories needs to be done with a lot of care [LRG20]. However, we will only consider tensor products with $𝕜\Bbbk$-linear categories, which are better behaved and allow our naive approach to work.
The restriction in the following definition to $m\geq 3$ is not necessary in general. However, as we will have naturally this assumption and it makes the formulas more concise we present this special case.
@@
$\left(x,m\right)\otimes\left(x^{\prime},m^{\prime}\right)$ $\mapsto\left(xx^{\prime},xm^{\prime}+mx^{\prime}\right)$
@@
$x_{1}\otimes...\otimes x_{m}$ $\mapsto\eta\left(x_{1},...,x_{m}\right).$
@@ 2.2. Equivariant objects over categories
In this section we recall the basic notions of equivariant objects and how to interpret them as functors from a small category. We will apply this in the next section to compute the equivariant characteristic morphisms of these objects.
@@
If there is a $\Gamma$-action on $T$ the characteristic morphism from Definition 2.8 admits an equivariant incarnation keeping track of the $\Gamma$-action.
@@
Let $\mathcal{X}$ be a $𝕜\Bbbk$-linear category, $\mathcal{M}$ a $𝕜\Bbbk$-central bimodule and let $\mathcal{G}$ be a $\Gamma$-equivariant $\mathcal{X}$-module. Then _$\_\otimes\mathcal{G}$ induces the $\Gamma$-equivariant characteristic morphism
@@
While $\mathrm{RHom}\left(\_,\mathcal{G}\right)$ induces the $\Gamma$-equivariant dual characteristic morphism
@@
Furthermore we can make a $\Gamma$-equivariant object $X$ into a functor via fixing the image of * to be $X$ and the structure morphism given by the map on morphism spaces. This gives us two inverse constructions.
@@
We in particular use the following notation for the category of equivariant objects over $\mathcal{I}$ in $\mathcal{C}$ by
@@ 3. Equivariant characteristic morphism of Injectives
We study the characteristic morphism introduced in the previous chapter for the subcategory of injectives. We start by discussing how we can reinterpret Hochschild cohomology and the characteristic morphism by the work of [Toe07]. We then use these insights to prove that we can treat the full subcategory of injectives similarly to a tilting object.
@@ 3.1. Interpreting characteristic morphisms of a subcategory as pullback
We use observations from [Toe07] to study the characteristic morphism
@@
$c_{\Gamma,T}:\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\to\mathrm{Ext}_{\mathcal{C}_{\Gamma}}^{*}\left(T,\mathcal{M}\otimes_{\mathcal{C}}T\right).$
@@
Let $\mathcal{C}$ be an abelian category and $\mathcal{I}\subset\mathcal{C}$ a subcategory. Then denote by $\iota_{I}:\mathcal{I}\hookrightarrow\mathcal{C}\subset\mathcal{D}_{\mathrm{dg}}^{b}\left(\mathcal{C}\right)$ the inclusion of $\mathcal{I}$.
@@
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\Gamma,\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota_{T},\psi_{\mathcal{M}}\circ\iota_{\mathcal{G}}\right)\cong\mathrm{Ext}_{\mathcal{C}_{\Gamma}}\left(T,T\otimes\mathcal{M}\right).$$c_{\Gamma,T}\cong\iota_{T,*}$
In particular we can use this interpretation to compute the equivariant characteristic morphism of the subcategory $\mathcal{I}\subset\mathcal{C}$:
$\mathrm{HH}^{*}\left(\mathcal{C},\mathcal{M}\right)\cong\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right),\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\mathrm{Id},\psi_{\mathcal{M}}\right)$$\mathrm{Ho}\left(\mathrm{dg}\mathrm{Cat}\right)\left(\mathcal{I},\mathcal{D}^{b}_{\mathrm{dg}}\left(\mathcal{C}\right)\right)\left(\iota,\psi_{M}\circ\iota\right)\cong\mathrm{Ext}^{*}_{\mathcal{C}_{\mathcal{I}}}\left(\iota,\psi_{M}\circ\iota\right).$$c_{\iota_{\mathcal{I}},\mathcal{I}}=\iota_{\mathcal{I},*}$
The above observation is in particular useful as explicit computations for bimodules between $𝕜\Bbbk$-linear categories, or, even more, $\mathcal{A}_{\infty}$/$\mathrm{dg}$-categories can be very unwieldy. However, the above observation allows us to compute the equivariant characteristic morphisms of the $\mathcal{I}$-$\mathcal{C}$ bimodule $\iota_{\mathcal{I}}$ in special cases.